OpenStudy (anonymous):

A man divided $9,000 among his wife, son, and daughter. The wife received twice as much as the daughter, and the son received $1,000 more than the daughter. How much did each receive?

OpenStudy (shinalcantara):

let x be the amount the wife received y be the amount for the daughter z for the son Since the total amount is 9,000 then x + y + z =9,000 ----equation 1 From the statement "The wife received twice as much as the daughter" x = 2y -----equation 2 "the son received $1,000 more than the daughter" z = y +1000 -----equation 3 substitute equations 2 and 3 to equation 1 x + y + z =9,000 2y + y + y + 1000 = 9,000 4y = 9,000 - 1,000 4y = 8000 y = 2,000 substitute to equations 2 and 3 x = 2y x = 2(2,000) x = 4,000 z = y + 1000 z = 2,000 + 1,000 z = 3,000